OpenStudy (anonymous):

If you know that two legs of a right triangle are 10 and 11 units long, what is the triangle's area? If necessary, round your answer to two decimal places.

OpenStudy (anonymous):

\[A=\frac{ ab}{ 2 }\]

OpenStudy (anonymous):

\[\frac{1}{2}\times 10\times 11\]

OpenStudy (dan815):

:O its satellite

OpenStudy (anonymous):

\[A=\frac{ 10*11 }{ 2 }\]

OpenStudy (anonymous):

\[A=\frac{ 110 }{ 2 }\]

OpenStudy (anonymous):

\[A=55\]

OpenStudy (anonymous):

55?

OpenStudy (anonymous):

@jessicaf correct

OpenStudy (anonymous):

thats all? kay thanks :)

OpenStudy (anonymous):

@jessicaf no problem